# Multitask Learning Can Improve Worst-Group Outcomes

### 05 Dec 2023 v1

@@ 5.4 Impact of Pre-Training
Finetuning pre-trained models is arguably the de-facto paradigm in machine learning (Devlin et al., 2018; Dosovitskiy et al., 2020; Dery et al., 2021b). Consequently, our experiments so far have exclusively focused on pre-trained models. In this section, we wish to understand the effect of deviating from this paradigm on our MTL approach. We thus compare against JTT and ERM when the model is trained from scratch instead of starting with a pre-trained model.
@@
Tables 5 and 6 depict our results on Waterbirds and Civilcomments-small, respectively. Our results show that pre-training is critical for setting up regularized MTL as a viable remedy against poor worst-group outcomes. We posit the following explanation for this outcome. Note that our informal motivation in Section 2 presupposes an ability to solve the auxiliary task to a reasonable degree. Solving the MLM and MIM tasks effectively from scratch with only the inputs of the relatively small supervised dataset is difficult. This poor performance on the auxiliary task translates to an inability to constrain the use of the spurious features on the end-task. Consistent with prior works (Tu et al., 2020; Wiles et al., 2022), our recommendation to practitioners is that our approach be used during finetuning of pre-trained models in order to be maximally effective.
@@
5.5 Going Beyond the Pre-training Objective
@@
Previous works in multitasking with self-supervised objectives suggests that different auxiliary objectives have disparate impact on end task performance (Dery et al., 2023). Out of curiousity about the impact of the choice of auxiliary objective, we explore the impact of going beyond the model’s original pre-training objective. For the Waterbirds dataset, we experiment with SimCLR – a constrastive prediction task based on determining whether two distinct augmented images originate from the same base image (Chen et al., 2020). For BERT experiments on Civilcomments-small, we substitute the standard masked language modeling (MLM) task with causal language modeling (CLM) as the auxiliary task. From the results in Table 7, we observe that SimCLR’s performance closely resembles that of the MIM pre-training objective, whereas CLM shows relatively inferior results compared to MLM. We hypothesize that BERT’s intrinsic bidirectional attention mechanism and non-autoregressive nature are not ideally suited for causal language modeling (Song et al., 2019), resulting in the model underperforming in our multitask setup. Given the sensitivity of model performance to the choice of the replacement objective, we proffer a practical recommendation to practitioners: use the pre-training objective as the auxiliary task. This is in line with recent work on best practices for fine-tuning pre-trained models (Goyal et al., 2023).
6 Related Work
Multitask Learning
Multitask learning is a common way for ML practitioners to improve the average performance of their models (Ruder, 2017; Ruder et al., 2019; Liu et al., 2019). Whilst work like Mao et al. (2020) has show that multitasking can improve the adversarial robustness of models, the impact of multitasking on worst-group outcomes has been relatively unexplored. Our work is inspired by Gururangan et al. (2020) who introduce constructing auxiliary objectives directly from end-task data for continued pre-training (they dub this Task Adaptive Pre-training – TAPT). Following Dery et al. (2021b; 2023), we multitask this auxiliary task with the end-task but unlike these works, our focus is on improving the worst-case group accuracy of the final model. Michel et al. (2021) consider balancing worst and average performance in multitask learning. In their work, they consider a set of equally important end tasks, all of which they want their model to perform well with respect to. In our work, we consider the asymmetrical multitask setting where the auxiliary task is only present in as much as it helps improves our target metric on the end task.
Robustness using group demographics.
Our multitask learning approaches is primarily designed for settings with limited-to-no group annotations. However, many DRO approaches assume the presence of group annotations for all training points. Among the approaches that leverage group information, Group Distributionally Robust Optimization Sagawa et al. (2020a) is the most popular technique that tries to minimize the maximum loss over the sub-groups. Goel et al. (2021) presented Model Patching, a data augmentation method designed to enhance the representation of minority groups. FISH, proposed by Shi et al. (2022), focuses on domain generalization via inter-domain gradient matching. In the settings where group annotations are expensive (financially or in terms of man-power) to procure, these methods are not viable options.
Robustness without group demographics.
Extensive research has been dedicated to addressing the challenges of worst-group generalization in the more realistic scenario where access to group annotations during training are not available. GEORGE Sohoni et al. (2020) adopts a clustering-based methodology to unveil latent groups within the dataset and subsequently employs groupDRO for improved robustness. Learning from Failure (LfF) Nam et al. (2020) introduces a two-stage strategy. In the first stage, an intentionally biased model aims to identify minority instances where spurious correlations do not apply. In the second stage, the identified examples are given increased weight during the training of a second model. Just Train Twice (JTT) method Liu et al. (2021) follows a similar principle by training a model that minimizes loss over a reweighted dataset. This dataset is constructed by up-weighting training examples that are misclassified during the initial few epochs.
Our regularized MTL approach has several advantages over these methods, even though they are all deployed in the same limited-to-no group annotations settings. As we have demonstrated, our approach is capable of improving both worst-group and average performance unlike the other approaches that are targeted against worst-group error only. Secondly, due to the already widespread usage of multitask learning by many ML practitioners, implementing our modification represents minimal overhead as opposed to introducing one of the above bespoke approaches.
7 Conclusion
We presented an empirical investigation of the impact of multitasking on worst-case group accuracy. We found that deploying multitasking, as is, doest not consistently improve upon worst-group outcomes. We have shown that whilst DRO methods, like JTT, display superior performance when models are trained from scratch, this is not the case in the currently more widespread setting of fine-tuning a pre-trained model. Specifically when fine-tuning, our method – regularized multitasking of the end-task with the pre-training objective constructed over end task data – leads to improvements in worst-case group accuracy over JTT. Our work has demonstrated that it is possible to design a single, simple method which improves both worst-case group accuracy and average accuracy regardless of the availability of group annotations. Since multitask learning is already a standard part of many practitioner’s tool box, and our modification to adapt it against worst-group accuracy is simple, our approach requires minimal overhead to integrate into existing systems compared to bespoke DRO approaches. We thus encourage practitioners to introduce our modification to their MTL pipelines as an essentially free way of improving worst-group performance without sacrificing gains in average performance.
In order to keep an apples-to-apples comparison with DRO approaches, we have primarily focused on multitasking with auxiliary objectives based on end task data only. For future work, it would be interesting to explore more deeply the impact of multitasking with auxiliary objectives based on external data. It would also be interesting to leverage meta-learning to dynamically adapt the auxiliary tasks towards improving worst case group outcomes (Dery et al., 2021b; 2023).
8 Acknowledgements
This work was also supported in part by the Tang AI Innovation Fund, Defence Science and Technology Agency Singapore, National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, and funding from Meta, Morgan Stanley, Amazon, Google, Schmidt Early Career Fellowship and Apple. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies.
Appendix A Appendix
A.1 Training Details
For $\mathbf{T}_{\mathrm{prim}}$, we employ a task-specific classification head of a single-layer multi-layer perceptron (MLP). For $\mathbf{T}_{\mathrm{aux}}$ we leverage the pre-trained MLM and MIM heads from BERT and ViT, respectively. We utilize the embedding of the [CLS] token from the base model through this MLP for classification. To facilitate effective multitask training, we adopt a task-heterogeneous batching scheme Aghajanyan et al. (2021). This facilitates the accumulation of gradients across tasks prior to each parameter update, contributing to improved training efficiency and convergence. Lastly, to ensure proper scaling, the L1 loss is normalized by the number of parameters in the shared representation.
@@ A.1 Training Details
A.2 Bayes optimal model for dimension-independent reconstruction under noised inputs

### 28 Feb 2024 v2

@@ 5.4 Impact of Pre-Training
Fine-tuning pre-trained models is arguably the de-facto paradigm in machine learning (Devlin et al., 2018; Dosovitskiy et al., 2020; Dery et al., 2021b). Consequently, our experiments so far have exclusively focused on pre-trained models. In this section, we wish to understand the effect of deviating from this paradigm on our MTL approach. Thus, we compare against JTT and ERM when the model is trained from scratch instead of starting with a pre-trained model. Tables 5 and 6 depict our results on Waterbirds and Civilcomments-small, respectively. Our results show that pre-training is critical for setting up regularized MTL as a viable remedy against poor worst-group outcomes. We posit the following explanation for this outcome. Note that our informal motivation in Section 2 presupposes an ability to solve the auxiliary task to a reasonable degree. Solving the MLM and MIM tasks effectively from scratch with only the inputs of the relatively small supervised dataset is difficult. This poor performance on the auxiliary task translates to an inability to constrain the use of the spurious features on the end-task. Consistent with prior works (Tu et al., 2020; Wiles et al., 2022), our recommendation to practitioners is to use our approach during the fine-tuning of pre-trained models to be maximally effective.
@@
6 Related Work
  • Multitask Learning. Multitask learning is a common stratergy for ML practitioners to improve the average performance of their models (Ruder, 2017; Ruder et al., 2019; Liu et al., 2019). While work like Hendrycks et al. (2019; 2020); Mao et al. (2020) have shown that multitasking can improve the adversarial and out-of-distribution robustness of models, the impact of multitasking on worst-group outcomes has been relatively unexplored. Makino et al. (2022) propose generative multitask learning (GMTL), a method that bolsters robustness to target shift by conditioning the input on all available targets, thereby addressing challenges associated with target-causing confounders and spurious dependencies between input and targets. However, it is important to note that their approach necessitates all target annotations during training, a requirement we do not assume in our scenario. Our work is inspired by Gururangan et al. (2020), who introduce constructing auxiliary objectives directly from end-task data for continued pre-training (they dub this Task Adaptive Pre-training – TAPT). Following Dery et al. (2021b; 2023), we multitask this auxiliary task with the end-task. However, unlike these studies, our focus is on improving the worst-case group accuracy of the final model. The work by Michel et al. (2021) explores the balancing of worst and average performance in multitask learning. In their study, they focus on a set of equally important end tasks, striving for proficient model performance across all of them. In contrast, our work delves into the asymmetrical multitask setting, where the presence of the auxiliary task is determined by its contribution to enhancing our target metric in the end task.
  • Robustness using group demographics. Our multitask learning approach is primarily designed for settings with limited-to-no group annotations. However, many DRO approaches assume the presence of group annotations for all training points. Among the approaches that leverage group information, Group Distributionally Robust Optimization Sagawa et al. (2020a) is the most popular technique that tries to minimize the maximum loss over the sub-groups. Goel et al. (2021) presented Model Patching, a data augmentation method designed to enhance the representation of minority groups. FISH, proposed by Shi et al. (2022), focuses on domain generalization via inter-domain gradient matching. In the settings where group annotations are expensive (financially or in terms of human resources) to procure, these methods are not viable options.
  • Robustness without group demographics. Extensive research has been dedicated to addressing the challenges of worst-group generalization in the more realistic scenario where access to group annotations during training is unavailable. GEORGE Sohoni et al. (2020) adopts a clustering-based methodology to unveil latent groups within the dataset and subsequently employs groupDRO for improved robustness. Learning from Failure (LfF) Nam et al. (2020) introduces a two-stage strategy. In the first stage, an intentionally biased model aims to identify minority instances where spurious correlations do not apply. In the second stage, the identified examples are given increased weight during the training of a second model. Just Train Twice (JTT) method Liu et al. (2021) follows a similar principle by training a model that minimizes loss over a reweighted dataset. This dataset is constructed by up-weighting training examples misclassified during the initial few epochs. Our regularized MTL approach has several advantages over these methods, even though they are all deployed in the same limited-to-no group annotations settings. As we have demonstrated, our approach can improve both worst-group and average performance, unlike the other approaches targeted against worst-group error only. Secondly, due to the widespread usage of multitask learning by many ML practitioners, implementing our modification represents minimal overhead instead of introducing one of the above bespoke approaches.
7 Conclusion
In this work, we presented an empirical investigation of the impact of multitasking on worst-group outcomes. We found that deploying multitasking, as is, does not consistently improve upon worst performing groups. We have shown that while DRO methods, like JTT, display superior performance when models are trained from scratch, this is not the case in the currently more widespread setting of fine-tuning a pre-trained model. Specifically, when fine-tuning, our method – regularized multitasking of the end-task with the pre-training objective constructed over end-task data – leads to improvements in worst-case group accuracy over JTT. Our work has demonstrated that it is possible to design a single, simple method that improves both worst-case group accuracy average accuracy regardless of the availability of group annotations. Since multitask learning is already a standard part of many practitioner’s toolbox, and our modification to adapt it against worst-group accuracy is simple, our approach requires minimal overhead to integrate into existing systems compared to bespoke DRO approaches. We thus encourage practitioners to introduce our modification to their MTL pipelines as an essentially free way of improving worst-group performance without sacrificing gains in average performance.
In order to keep an apples-to-apples comparison with DRO approaches, we have primarily focused on multitasking with auxiliary objectives based on end-task data only. For future work, it would be interesting to explore the impact of multitasking with auxiliary objectives based on external data more deeply. It would also be interesting to leverage meta-learning to dynamically adapt the auxiliary tasks towards improving worst-case group outcomes (Dery et al., 2021b; 2023). Additionally, we leave the study on the generalizability of our method to adversarial robustness, domain shift, and label shift to future work.
8 Acknowledgements
This work was also supported in part by the Tang AI Innovation Fund, Defence Science and Technology Agency Singapore, National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, and funding from Meta, Morgan Stanley, Amazon, Google, Schmidt Early Career Fellowship and Apple. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies.
Appendix A Appendix
A.1 Training Details
For $\mathbf{T}_{\mathrm{prim}}$, we employ a task-specific classification head of a single-layer multi-layer perceptron (MLP). For $\mathbf{T}_{\mathrm{aux}}$, we leverage the pre-trained MLM and MIM heads from BERT and ViT, respectively. We utilize the embedding of the [CLS] token from the base model through this MLP for classification. To facilitate effective multitask training, we adopt a task-heterogeneous batching scheme Aghajanyan et al. (2021). This facilitates the accumulation of gradients across tasks prior to each parameter update, contributing to improved training efficiency and convergence. Lastly, to ensure proper scaling, the L1 loss is normalized by the number of parameters in the shared representation.
A.2 Going Beyond the Pre-training Objective
@@
Previous works in multitasking with self-supervised objectives suggest that different auxiliary objectives have disparate impacts on end task performance (Dery et al., 2023). Out of curiosity about the impact of the choice of auxiliary objective, we explore the impact of going beyond the model’s original pre-training objective. For the Waterbirds dataset, we experiment with SimCLR – a constrastive prediction task based on determining whether two distinct augmented images originate from the same base image (Chen et al., 2020). For BERT experiments on Civilcomments-small, we substitute the standard masked language modeling (MLM) task with causal language modeling (CLM) as the auxiliary task. From the results in Table 7, we observe that SimCLR’s performance closely resembles that of the MIM pre-training objective, whereas CLM shows relatively inferior results compared to MLM. We hypothesize that BERT’s intrinsic bidirectional attention mechanism and non-autoregressive nature are not ideally suited for causal language modeling (Song et al., 2019), resulting in the model underperforming in our multitask setup. Given the model performance’s sensitivity to the replacement objective’s choice, we proffer a practical recommendation to practitioners: use the pre-training objective as the auxiliary task. This aligns with recent work on best practices for fine-tuning pre-trained models (Goyal et al., 2023).
@@
A.3 Bayes optimal model for dimension-independent reconstruction under noised inputs
@@
Appendix B Broader Impact Statement
In terms of broader impact, our work has beneficial implications for group fairness and mitigating demographic-based bias in machine learning systems. Specifically, we present a simple approach to improving worst-case group error. This means that our work contributes positively to certain groups that would otherwise be impacted by the poor performance of ML models. We have also provided a new lens from which to view the problem of the impacts of multitasking, resulting in showing that it is tool able against poor group-based outcomes. However, our method introduces compute overhead due to the optimization of multiple objectives. This results in increased power consumption and, thus, greenhouse emissions during the training of models. Nevertheless, introducing our simple regularization modification to existing MTL implementations represents a negligible technical overhead than introducing JTT or BR-DRO to target worst-group error.
Table 8: Sensitivity analysis of all the methods to different hyperparameters wrt worst-group accuracy.
Dataset Method Seed Learning Rate Batch Size Up T $\lambda_{aux}$ $\lambda_{reg}$
Waterbirds ERM $0.0734$ $0.5717$ $-0.4384$- -- -- --
JTT $0.0573$- $-0.380$ $0.0246$ $0.1862$- --
ERM + MT + L1 $-0.0314$- $0.0142$- -- $0.0943$ $0.4114$
groupDRO .$-.1688$ $0.196$ $-0.1384$- -- -- --
Civilcomments-Small ERM $0.0617$ $-0.8$ $0.1029$- -- -- --
JTT $-0.0315$- $0.2104$ $-0.0047$ $0.1111$- --
ERM + MT + L1 $0.0677$- $0.0453$- -- $0.0601$ $-0.2955$
groupDRO $-0.0392$ $-0.7352$ $-0.1726$- -- -- --
Sensitivity to Hyperparameters:
For evaluating the sensitivity of worst group accuracy to various hyperparameters across different methods, we employ Kendall’s rank coefficient $\tau$ Kendall (1938). To adhere to our computation budget, we opt for the optimal learning rate utilized in ERM for both JTT and ERM+MT+L1. Our sensitivity analysis from Table 8 reveals distinct preferences in learning rates. The ViT model applied to waterbirds, demonstrating a preference for higher learning rates. Conversely, when trained on the civilomments-small dataset, BERT exhibits a preference for lower learning rates. The epoch at upsampling is also a sensitive hyperparameter for JTT. The sensitivity to different seed values is negligible across all the methods. Notably, our method displays the most minor sensitivity to changes in batch size. However, it is noteworthy that our method is most responsive to variations in the regularization parameter ($\lambda_{reg}$), with lower values resulting in higher worst group accuracy.